A mass of 2.3 kilograms is placed on a horizontal frictionless surface against an uncompressed spring with spring constant 1158.7 N/m. The mass is pushed against the spring until the spring is compressed a distance 0.76 m and then released. How high (vertically) in m does the mass rise from the original height before it stops (momentarily).
Answer #1

Energy, E = 0.5 kx^2

E = 0.5 x 1158.7 x 0.76^2 = 334.63 J

Also,

Energy, E = mgh

Height, h = 334.63/2.3 x 9.8 = 14.85 m

Comment in case any doubt please rate my answer....